OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 4x + 3y = 10 4x + 5y = 14

Parth (parthkohli):

First subtract both equations, find the value for y, then find x.

Parth (parthkohli):

Can you manage now?

Parth (parthkohli):

Subtract the first equation from the second for easier work.

Parth (parthkohli):

\(\Rightarrow (4x - 4x) + (5y - 3y) = 14 - 10 \)

Parth (parthkohli):

\(\Rightarrow 2y = 4 \) \(\Rightarrow y = 2 \)

Parth (parthkohli):

As we already know that y = 2 we could do a little trick here. \(\Rightarrow 4x + 3(2) = 10 \)

OpenStudy (anonymous):

so what would x equal? That's what I don't understand, but thank you!